What level of income is required for a household to be in the top 10% of households?

Suppose household income (X) in the United States is distributed normally with an average realization of $50,000 and standard deviation $25,000.


1. What level of income is required for a household to be in the top 10% of households?
Answer #1

data:

x-bar.png=$50,000

  σ =25000

    x     =?

    z=10%

solution:

z                = (x-μ)/σ,

z× σ,        = (x-μ)

z× σ+μ     =x

0.1 ×25000+50000=x

x                =52500




answered by: Sumayya hasan Khan